Integrate the following by substitution method :
1/(ex +e-x)2

Grade:12th pass

1 Answers

Aditya Gupta
5 years ago
hello avni. this is an easy ques.
1/(ex +e-x)= e^2x/(e^2x+1)^2 
now substitute e^2x + 1 = y, and the ques becomes as easy as that!
- 1/2(e^2x + 1) + C is answer